dy/dx+ycotx=5ecosx

1
Expert's answer
2022-02-22T22:43:41-0500

Integrating factor


"\mu(x)=e^{\int\cot xdx}=e^{\ln (\sin x)}=\sin x"

"\int \cot xdx=\int\dfrac{\cos x}{\sin x}dx=\ln(|\sin x|)+C"

"y'\sin x+y\cos x=5e^{\cos x}\sin x"

"d(y\sin x)=5e^{\cos x}\sin xdx"

Integrate


"\int d(y\sin x)=\int 5e^{\cos x}\sin xdx"

"y\sin x=-5e^{\cos x}+C"

"y=-5e^{\cos x}\csc x+C\csc x"